VEGAS Team Meeting

Attendees: Ray, Joe, Jason, Melinda, Mark, John, Richard, Paul, Dan, (Steve)

Agenda:

  1. vegasdm: measpwr in pol 2 not working?
  2. data normalization when exposure == 0
  3. ADC Calibration (should be complete)
  4. Review of “mode 1 ripple” fix.
  5. Status of “1ms integration time” problem (should have been 0.5 ms).
  6. Status of blanking time / mixed mode discussion and implementation.
  7. Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues
  8. Status of shared-risk astronomical observations: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerAstronomicalObservations
  9. Plans for testing for this week.

Meeting notes:

  1. vegasdm: measpwr in pol 2 not working?
    • need to determine if power is being measured in dBm consistently throughout at snap blocks.
    • similarly looking at measpwr last Friday at 1st ADC snap block and there were -21to -31 dBm variations.
    • need some log details from samplers to obtain data for analysis
  2. data normalization when exposure == 0
    • Part one: "6/8 problem" - Obs. with banks with 8 subbands and some with 6 subbands did not fill fits data properly, Bob determined 8 banks are being written but one banks had 6 banks that wrote eight - looks to be in manager; config 'knows about the bank with 6 subbands. A method to fill the data correctly later would save obs. Work around exists. Problem needs to be fixed in Manager/Config though
    • Part two:good news: according to my unit tests, the manager does not simply crash when encountering exposure's of zero and normalizing with this (dividing by zero). SDFits will deal with this currently. If integration time is zero you 'NaN' will flag the bad floating point value
  3. ADC Calibration:
    • calibration code
    • loading calibration results
    • recalibration of all the ADCs needed due to bug in calibration code (see below)
    • create a standalone python code module so that this procedure can be passed off to DIBAS
  4. Review of “mode 1 ripple” fix.
    • Amanda took a look at the mode 1 ripple test data that Richard took last week. Compared to the previous BOF file, most of these spectra were actually worse the previous data -- i.e., larger ripples and ripples in both polarizations. Details of the observing: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometer2014Apr15 . I've attached PDFs of ifnum=0 as representative sample. These should be compared to the beforeADC_mode1 plots. The only exceptions were ifnum=2, plnum=1 and ifnum=3, plnum=0.
    • Glenn needs to look at this in consultation with Amanda
  5. Status of “1ms integration time” problem (should have been 0.5 ms).
    • Proposed solution is to change computation to simply 'round' and not 'round up' [agreed]
    • Noted that granularity is poor at low exposures but we should leave it as is for now to keep speeds high (Ray)
  6. Status of blanking time / mixed mode discussion and implementation.
    • [needs Joe who is absent from meeting]; Manager updates will be needed as part of this fix.
  7. Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues
  8. Status of shared-risk astronomical observations: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerAstronomicalObservations
  9. Plans for testing for this week.
    • Thursday - obtain some log details from samplers to obtain data for analysis
    • Thursday - recalibration of all the ADCs needed due to bug in calibration code during receiver checkout (Jason)
    • Thursday - release code for rounding (Jason)
    • Friday -

Other

  1. Please make sure to properly code your timesheets for the current time period and going forward.

- Software engineers should be using a GBT ops software CM&E account

Account 893252

- Electronic engineers should be using a GBT ops electronics CM&E account

Account 432510

- Commissioning activities should be getting charged to hardware support and testing (432510)

Account 432510

- scientists assisting with shared-risk commissioning should be charging an observer support account.

432540 for the most part. Accounts 142023 (GB-based staff) and 142013 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)

-- MartyBloss - 2014-04-14
Topic revision: r3 - 2014-04-21, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback